OpenStudy (anonymous):

Part 1: If y=arcsin x, find cos y and tan y. Part 2: If y=arctan x, find cos y and sin y. **How do you start this problem off? Thanks a bunch!! :)

OpenStudy (agent0smith):

Draw a right triangle using this, make one angle y y=arcsin x, then siny = x or siny = x/1

OpenStudy (agent0smith):

Then find cosy and tany from the triangle. Same thing for part 2

OpenStudy (anonymous):

okay, so i draw the triangle like this right? (attached) what do i do next? :/ not sure how to do the arcsin x of it :(

OpenStudy (anonymous):

oh whoops it attached sideways haha :P

OpenStudy (agent0smith):

y=arcsin x, then siny = x or siny = x/1

OpenStudy (anonymous):

okay.. did i do it right? :/

OpenStudy (agent0smith):

Well that side should just be x. Now find the remaining side.

OpenStudy (anonymous):

okay so should it be x and then 1 and would the remaining side be (sq.rt. 1- x^2) ?

OpenStudy (agent0smith):

Yep now find cosy and tany.

OpenStudy (anonymous):

how do i do that? :/

OpenStudy (agent0smith):

You know how to do it if you've gotten this far ;)

OpenStudy (anonymous):

okay :P lemme try to remember hehe one sec :)

OpenStudy (anonymous):

cos(y)=sq.rt 1-x^2

OpenStudy (anonymous):

? :/

OpenStudy (agent0smith):

Now tany.

OpenStudy (anonymous):

tan y = 1 ?

OpenStudy (anonymous):

is that right? :/ or did i mix it up lol

OpenStudy (agent0smith):

Tan is opposite/adjacent...

OpenStudy (anonymous):

oh so x/sq.rt 1-x^2 ?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

awesome!! :) so for part 2, would it be this? cos y= (sq.rt 1-x^2)/x and would sin y =sq.rt 1-x^2 ??

OpenStudy (agent0smith):

idk... draw me a triangle :P

OpenStudy (agent0smith):

No it won't be tany = x/1, then use pythag.

OpenStudy (anonymous):

okay... are these correct for part 2? :)

OpenStudy (anonymous):

sorry it's sideways again haha... don't know why it's attaching like that :/

OpenStudy (agent0smith):

You're using pythag wrong :P

OpenStudy (anonymous):

oh oops... lemme redraw it haha one sec :P

OpenStudy (agent0smith):

Just change all the - to + and you're done!

OpenStudy (anonymous):

is this better?

OpenStudy (anonymous):

oh shoot, i forgot the sq.rt :/ so it should be sq.rt x^2 + 1 is the hypotenuse and tan y=x/1 cos y= 1/ sq.rt x^2 + 1 sin y = x/ sq.rt x^2 + 1 is that right?

OpenStudy (agent0smith):

Yep

OpenStudy (anonymous):

ah! awesome!! thanks so much!!! :)

OpenStudy (agent0smith):

Welcome :)